# Competitive Online Transportation Simplified

### v1

@@ 1 Introduction
Let us start by defining the problem:
@@ 1.1 The Story to Date
In the mid 1990’s, one of the three natural followup research directions/questions, proposed in [11], was to determine the the optimal deterministic competitive ratio for the online transportation problem. In fact it was conjectured that:
@@ 1.2 Our Contribution
The key insight in our analysis of algorithm $\mathcal{A}$ is to note that that the final parking spots of the cars in algorithm $\mathcal{A}$ essentially matches the optimal bottleneck matching. This observation is a relatively straightforward consequence of the properties of the algorithm. We then extend algorithm $\mathcal{A}$ for a power-of-two tree metric $T$ to an algorithm $\mathcal{B}$ for a general metric space $M$ as in [8]. We conclude Section 2 by noting that a straightforward consequence of this observation is that algorithm $\mathcal{B}$ is (8m7)(8m-7)-competitive for the itinerant car problem.
@@ 2.1 Algorithm $\mathcal{A}$ for a Power-of-Two Tree Metric
We start with some definitions. Let $T=(V,E)$ be a power-of-two tree metric. That is, one garage is located at each vertex in $V$, each edge in $E$ has an associated weight that is of the form $2^{j}$ for some integer $j\in[0,n]$, and the distance between any two points is the length of the unique path between these points in the tree $T$. A garage is full if the number of cars it contains equals its capacity, and unfull otherwise. Define $T_{j}(i)$ to be the level $j$ tree for garage $i\in[m]$, that is the portion of $T$ reachable from garage $i$ without traversing any edges with weight $2^{j}$ or greater. Define $W_{j}(i)$ as an arbitrary depth-first-search walk/tour on $T_{j}(i)$, starting from $i$, which we will call the level $j$ walk for garage $i$. Alternatively $W_{j}(i)$ can be thought of as an Eulerian tour that traverses each edge in $T_{j}(i)$ twice, once in each direction. We now give an algorithm for the itinerant car problem on a power-of-two tree metric.
@@
Each car $t$ that arrives at garage $i$ first drives along the route of the walk $W_{0}(i)$, then along the route of the walk $W_{1}(i)$, then along the route of $W_{2}(i)$, etc. If while traversing $W_{j}(i)$ this car $t$ arrives at a garage $h$, the following occurs:
@@ 2.2 Analysis of Algorithm $\mathcal{A}$
Line (1) follows since car $t$ arriving at garage $i$ will traverse walk $W_{j+1}(i)$ at most $(\ell(t)-j)$ times for $j\leq\ell(t)-1$, and an edge of weight $2^{j}$ will only appear in trees $T_{j+1}(i),T_{j+2}(i),\dots,T_{\ell(t)}(i)$. Line (5) follows since tree $T$ has m1m-1 edges. Line (7) follows from Lemma 3. ∎
@@ 2.3 Algorithm $\mathcal{B}$ and Analysis for General Metric Spaces
Inequality (11) follows from fact that the aggregate distance from the arrival location of each car in $M$ to the nearest garage is a lower bound to the optimal cost $\mbox{Opt}(I,M)$, and the fact that the distance between two locations in $T$ is an upper bound for the distance between these locations in $M$. Inequality (12) follows from Theorem 4. Inequality (13) follows from Lemma 5. ∎

### v2

@@ 1 Introduction
Let us start by defining the problem.
@@ 1.1 The Story to Date
In the mid 1990’s, one of the three natural followup research directions/questions, proposed in [11], was to determine the optimal deterministic competitive ratio for the online transportation problem. In fact it was conjectured that:
@@ 1.2 Our Contribution
The key insight in our analysis of algorithm $\mathcal{A}$ is to note that that the final parking spots of the cars in algorithm $\mathcal{A}$ essentially matches the optimal bottleneck matching. An optimal bottleneck matching of cars to parking spots in a tree $T$ is a min-cost matching, where the cost of assigning a car at vertex $u$ to a parking spot at vertex $v$ is the weight of the largest edge on the uvu-v path in $T$. This observation is a relatively straightforward consequence of the properties of the algorithm. We then extend algorithm $\mathcal{A}$ for a power-of-two tree metric $T$ to an algorithm $\mathcal{B}$ for a general metric space $M$ as in [8]. We conclude Section 2 by noting that a straightforward consequence of this observation is that algorithm $\mathcal{B}$ is (8m7)(8m-7)-competitive for the itinerant car problem.
@@ 2.1 Algorithm $\mathcal{A}$ for a Power-of-Two Tree Metric
We start with some definitions. Let $T=(V,E)$ be a power-of-two tree metric. That is, one garage is located at each vertex in $V$, each edge in $E$ has an associated weight that is of the form $2^{j}$ for some integer $j\in[0,n]$, and the distance between any two points is the length of the unique path between these points in the tree $T$. Further, cars will arrive only at vertices of $T$. A garage is full if the number of cars it contains equals its capacity, and unfull otherwise. Define $T_{j}(i)$ to be the level $j$ tree for garage $i\in[m]$, that is the portion of $T$ reachable from garage $i$ without traversing any edges with weight $2^{j}$ or greater. Note $T_{0}(i)=i$ and $T_{n+1}(i)=T$ for all $i\in[m]$. Define $W_{j}(i)$ as an arbitrary depth-first-search walk/tour on $T_{j}(i)$, starting from $i$, which we will call the level $j$ walk for garage $i$. Alternatively $W_{j}(i)$ can be thought of as an Eulerian tour that traverses each edge in $T_{j}(i)$ twice, once in each direction. We now give an algorithm for the itinerant car problem on a power-of-two tree metric.
@@
Each car $t$ that arrives at garage $i$ first drives along the route of the walk $W_{0}(i)$ (which is just the vertex at garage $i$), then along the route of the walk $W_{1}(i)$, then along the route of $W_{2}(i)$, etc. If while traversing $W_{j}(i)$ this car $t$ arrives at a garage $h$, the following occurs:
@@ 2.2 Analysis of Algorithm $\mathcal{A}$
Line (1) follows since car $t$ arriving at garage $i$ will traverse the edges in $T_{j+1}(i)$ at most $2(\ell(t)-j)$ times for $j\leq\ell(t)-1$, and an edge of weight $2^{j}$ will only appear in trees $T_{j+1}(i),T_{j+2}(i),\dots,T_{\ell(t)}(i)$. Line (5) follows since tree $T$ has m1m-1 edges. Line (7) follows from Lemma 3. ∎
@@ 2.3 Algorithm $\mathcal{B}$ and Analysis for General Metric Spaces
Inequality (11) follows from the fact that the aggregate distance from the arrival location of each car in $M$ to the nearest garage is a lower bound to the optimal cost $\mbox{Opt}(I,M)$, and the fact that the distance between two locations in $T$ is an upper bound for the distance between these locations in $M$. Inequality (12) follows from Theorem 4. Inequality (13) follows from Lemma 5. ∎